Assume that the banking system is loaned up and that any open-market purchase by the Fed directly increases reserves in the banks. If the required reserve ratio is 0.2, by how much could the money supply expand if the Fed purchased $2 billion worth of bonds?
Answer #1

Answer

Money supply = m*H

where H = Monetary Base = (C)Currency + (R)Reserves and m = Multiplier

Simple money multiplier (i.e. when Excess reserves = 0 and Currency to Deposit ratio = 0) = 1/rr

=> m = 1/rr where rr = required reserve ratio

Here rr = 0.2 and Here Reserves and hence H(Monetary Base increases by 2 billions because of Open Market purchase

here m = 1/0.2 = 5 and

Hence Money Supply will increase by $10 billion

Hence , the money supply could spend by $10 trillion if the Fed purchased $2 billion worth of bonds.